A cargo plane flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220 mph. The average speed on the way back was 200 mph. How many hours did the trip
there take? Show ALL work.

Answer by TimothyLamb(4379)

---
s = d / t
d = s * t
---
out:
d = 220 * (t - 1)
---
back:
d = 200 * t
---
200t = 220(t - 1)
200t = 220t - 220
220t - 200t - 220 = 0
20t = 220
t = 11
---
answer:
trip there took 10 hours, and the return trip took 11 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php